Riders on the Power Tower are launched skyward with an acceleration of 4g, after which they experience a period of free fall. What is a 60 kg rider's apparent weight during the launch?
Answer #1

The riders apparent weight during the launch is FN = mg + ma = (60) (9.8)-(60) (4g) = (60) (9.8)-(60) (4x88) = 2940 N